# A priori estimates for parabolic Monge-Ampère type equations

### v1

@@ 2.1. Step 1: $C^{2}$ estimates on the boundary
$\mathcal{L}\mathcal{T}u=0,\quad\ \text{for~{}the~{} linear~{} operator}\quad\mathcal{L}=:u_{t}u^{ij}\partial_{ij}-\partial_{t}.$
Thus, by (2.3), we have
$|\mathcal{L}\mathcal{T}(u-\phi)|=|\mathcal{L}\mathcal{T}\phi|=|u_{t}u^{ij}\partial_{ij}\mathcal{T}\phi-\partial_{t}\mathcal{T}\phi|\leq\underline{c}\Big{(}1+\sum_{i=1}^{n}u^{ii}\Big{)}$ (2.9)
for some positive constant $\underline{c}$ depending on $||\phi||_{\tilde{C}^{4}(\overline{Q}_{T})}$ and the positive lower and upper bounds for $u_{t}$. Moreover, on $\Omega\times\{0\}$, $\mathcal{T}(u-\phi)=0$ and
$|\mathcal{T}(u-\phi)|=\Big{|}-\sum_{\beta<n}B_{\alpha\beta}x_{n}(u_{\beta}-\phi_{\beta})\Big{|}\leq\overline{c}x_{n},\quad\text{on }\partial\Omega\times[0,T].$ (2.10)
@@
Then $\mathcal{L}w=-2au_{t}\sum_{i=1}^{n}u^{ii}$. Note that
$\sum_{i=1}^{n}u^{ii}\geq n(\prod_{i=1}^{n}u^{ii})^{1/n}\geq n(\det D^{2}u)^{-1/n}\geq nC^{-1/n},$
where $C$ depends on the maximum of $u_{t}$. If we take $a$ big enough, then
@@
$u_{t}(0,t)=\det D_{x^{\prime}x^{\prime}}u(0,t)u_{nn}(0,t)+P\big{(}D_{x^{\prime}x^{\prime}}u(0,t),D_{x^{\prime}x_{n}}u(0,t)\big{)},$
@@
$||D^{2}u||\leq C\quad\text{ on }~{}~{}\partial_{p}Q_{T}.$ (2.12)
Since $u_{t}$ is strictly positive and bounded, we get a lower bound for eigenvalues of $D^{2}u$ on the parabolic boundary. ∎
@@ 2.2. Step 2: global $C^{2}$ estimates
$-U_{t}\det D^{2}U=1\quad\text{ in }Q_{T}^{*}.$ (2.13)
Note that $Du(\Omega,t)$ is not necessarily convex, neither is $Q_{T}^{*}$. Here $Q_{T}^{*}$ is bounded and simply connected due to the smoothness of $u$ and formula (2.2). By estimates (2.3) and (2.4) above, we know that
$C^{-1}\leq-U_{t}\leq C\quad\text{and}\quad||D^{2}U||_{L^{\infty}(\partial_{p}Q_{T}^{*})}\leq C.$ (2.14)
@@
Differentiating the equation (2.13), we have
$\frac{U_{tk}}{U_{t}}+U^{ij}U_{ijk}=0\quad\text{ for any }k=1,\cdots,n,$
@@
$\frac{U_{tkl}}{U_{t}}-\frac{U_{tk}U_{tl}}{U_{t}^{2}}+U^{ij}U_{ijkl}-U_{ijk}U_{lpq}U^{ip}U^{qj}=0\quad\text{ for any }k,l=1,\cdots,n.$ (2.15)
If we define $\tilde{L}=\frac{1}{U_{t}}\partial_{t}+U^{ij}\partial_{ij}$, then $\tilde{L}DU=0$ and
$\tilde{L}U_{kk}$ $=\frac{U_{tkk}}{U_{t}}+U^{ij}U_{ijkk}=\frac{U_{tk}^{2}}{U_{t}^{2}}+U_{ijk}U_{pqk}U^{ip}U^{qj}$
by the identity (2.15). The last term of the above equality is nonnegative since $U$ is convex. Hence we have $\tilde{L}U_{kk}\geq 0$. Note that $U_{t}<0$ and $(U^{ij})\geq 0$, then by the maximum principle for linear parabolic equations, we obtain that
$\sup_{Q_{T}^{*}}U_{kk}=\sup_{\partial_{p}Q_{T}^{*}}U_{kk},$
@@
$1/C\leq\lambda(D^{2}u)\leq C\quad\text{ on }~{}~{}\overline{Q}_{T},$ (2.16)
where $C$ is a positive constant depending on $n,\Omega,c_{0},||\phi||_{\tilde{C}^{4}(\overline{Q}_{T})}$ and the uniform convexity of $\phi(\cdot,0)$.
@@ 2.3. Step 3: global $\tilde{C}^{2,\alpha}$ estimates
By regularity theory [30] for the equation (2.13), $U$ is smooth in the interior. Then by (2.14), we have the higher regularity for $u$ in $Q_{T}$.
@@
Recall that $u\in\tilde{C}^{4}(\overline{Q}_{T})$ is a parabolically convex solution to equation (2.1). Then, for any $Q\subset\subset Q_{T}$, we have
$||u||_{\tilde{C}^{4}(\overline{Q})}\leq C,$
where $C$ is a positive constant depending on $n,\Omega,c_{0},\phi$ and $dist(Q,Q_{T})$.
Let $h=u_{t}$, then $h$ satisfies
$h_{t}=a^{ij}h_{ij}.$
Here $a^{ij}=(\det D^{2}u)u^{ij}$ and $h$ are bounded from below and above by Theorem 2.4. By theories of linear parabolic equations [24], we can obtain the $\tilde{C}^{\alpha}$ estimate for $u_{t}$ on $\overline{Q}_{T}$.
@@
for some positive constants $C$ and $\alpha$ depending on $n,\Omega,c_{0},||\phi||_{\tilde{C}^{4}(\overline{Q}_{T})}$ and the uniform convexity of $\phi(\cdot,0)$.
@@
The interior higher regularity of $u$ can also be obtained by using the same method in Proposition 5.3 of [4] once we get the global $\tilde{C}^{2}$ estimates and the $\tilde{C}^{\alpha}$ estimate for $u_{t}$.
We split the proof of the global $\tilde{C}^{2,\alpha}$ estimate into two parts. A short-time existence theorem gives the estimate for $D^{2}u$ on $\Omega\times(0,2\epsilon]$. As for the estimate on the rest part of $Q_{T}$, considering the linear parabolic equation satisfied by tangential derivatives of $u$, we use the approach of Theorem 9.31 in [13] to deduce the Hölder regularity of the gradient, which is independent of the moduli of continuity of the principal coefficients.
Let $Y_{0}=(y_{0},t_{0})$ be any point on $\partial\Omega\times[\epsilon,T]$ with some constant $\epsilon>0$. Without loss of generality, assume that $y_{0}=0$, and on a small neighborhood of 00, $\partial\Omega$ is the image of $y_{n}=\rho(y^{\prime})$. Perform the transformation of coordinates $x=\Psi(y)$, and $\Psi:B_{\rho_{0}}(0,t_{0})\cap\overline{Q}_{T}\to\overline{\mathbb{R}_{n}^{+}}\times(0,+\infty)$, with
@@
with any $1\leq k<n$, where
$T_{k}=:\partial_{k}+\rho_{k}\partial_{n}\quad\text{ for }~{}k<n,\quad T_{n}=\partial_{n}.$
Take $G=B_{R_{0}}^{+}\times(t_{0}-R_{0}^{2},t_{0}]$, such that $G\subset\operatorname{Im}\Psi$. Then by direct computations,
$-\partial_{t}(T_{k}u)+u_{t}u^{ij}(T_{k}u)_{ij}$ $=u_{t}u^{ij}\rho_{kij}u_{n}+u_{t}u^{ij}\rho_{ki}u_{nj}+u_{t}u^{ij}\rho_{kj}u_{ni}$
$=:\overline{f}.$
Thus under $(x,t)$ coordinates,
$-\partial_{t}v+u_{t}u^{ij}P_{i}P_{j}v=Lv-u_{t}u^{ij}\rho_{ij}v_{n}+LT_{k}\phi.$
Here $L=-\partial_{t}+a^{ij}\partial_{ij}$, $P_{k}=\partial_{k}-\rho_{k}\partial_{n}$ for $1\leq k<n$, $P_{n}=\partial_{n}$ and
$a^{ij}=u_{t}u^{ij}+u_{t}u^{kl}\rho_{k}\rho_{l}\delta_{n}^{i}\delta_{n}^{j}-u_{t}u^{kj}\rho_{k}\delta_{n}^{i}-u_{t}u^{ik}\rho_{k}\delta_{n}^{j},$
where $\delta_{n}^{i}$ equals $1$ when $n=i$, otherwise 00, and the summation over $k,l$ is from $1$ to $n-1$. Let
$f=u_{t}u^{ij}\rho_{ij}T_{k}u_{n}-LT_{k}\phi+\overline{f}.$
@@
{$-v_{t}+a^{ij}v_{ij}&=f\quad\text{ in }G,\\ v&=0\quad\text{ on }F,\end{aligned}\right.$ (2.17)
@@
$\lambda I_{n}\leq(a^{ij})\leq\Lambda I_{n},$ (2.18)
and
$|f|\leq C_{0},\quad\quad|v|+|Dv|\leq C_{1}\quad\text{in}\quad\overline{G}$ (2.19)
for some positive constants $\lambda$, $\Lambda$, $C_{0}$ and $C_{1}$ depending on $n,\Omega,c_{0},||\phi||_{\tilde{C}^{4}(\overline{Q}_{T})}$ and the uniform convexity of $\phi(\cdot,0)$.
@@
We will use Lemma 2.9 to prove an oscillation estimate for $\frac{v}{x_{n}}$, which gives the Hölder estimate for $v_{n}$ on $F$, the flat boundary. Combining this with the interior estimates, Lemma 2.8, and with the help of Lemma 2.10, we will derive a global $\tilde{C}^{1,\alpha}$ estimate for $v$ in $\frac{1}{2}G:=B_{R_{0}/2}^{+}\times(t_{0}-R_{0}^{2}/4,t_{0}]$. Recall the definition of $v$, then the global Hölder estimate of $D^{2}u$ on $\Omega\times(\epsilon,T]$ follows from the equation.
@@
If we take $\sigma=r+\mu/\delta$, then $X,X_{1}^{\prime}\in Q_{\sigma,\delta}\subset Q_{R_{0}/2,\delta}$. Consider the function $v(X)-\ell\cdot x$, then by Lemma 2.8 and Lemma 2.10,
$\begin{split}|Dv(X_{1})-Dv(X_{1}^{\prime})|&\leq\sup_{Q(X_{1},\mu/2)}|Dv-\ell|\\ &\leq C\Big{(}\mu^{-1}\sup_{Q(X_{1},\mu)}|v(X)-\ell\cdot x|+\mu^{\alpha}\Big{)}.\end{split}$ (2.20)
@@
$\operatorname{osc}\limits_{Q_{\sigma,\delta}}\frac{v(X)}{x_{n}}\leq C\sigma^{\alpha}.$ (2.21)
@@
$\leq Cr^{\alpha}.$
Again by the claim, $|Dv(X_{1}^{\prime})-Dv(0,t_{0})|\leq C\sigma^{\alpha}\leq Cr^{\alpha}$, then we have proved the theorem by the triangle inequality. For the rest part of this proof, we will show the claim (2.21).
@@
$\inf_{F_{R,\delta}}\frac{v}{x_{n}}\leq 10\Big{(}\inf_{Q_{R/2,\delta}}\frac{v}{x_{n}}+R\sup|f|\Big{)}\quad\text{ for $R<1/4R_{0}$.}$ (2.22)
@@
Hence, $Lw\geq\sup|f|\geq Lv$ in $Q_{1,\delta}$, when $\delta$ is small enough. On $\{x_{n}=0\}\cap\partial_{p}Q_{1,\delta}$, $w=0=v$. On $\{x_{n}=\delta\}\cap\partial_{p}Q_{1,\delta}$, $w\leq\delta(1-|x^{\prime}|^{2})\leq\delta\leq v$ by (2.23). Additionally, on $\{|x^{\prime}|=1\}\cap\partial_{p}Q_{1,\delta}$ and $\{t=t_{0}-1\}\cap\partial_{p}Q_{1,\delta}$, $w<0\leq v$. Thus by maximal principle, we have $w\leq v$ in $Q_{1,\delta}$. Therefore, in $Q_{1/2,\delta}$,
@@
when $\delta$ is sufficiently small. Now without the assumption (2.23), let
@@
$\begin{split}\Big{(}R^{-n-2}\int_{Q^{\prime}_{R,\delta}}(Mx_{n}-v)^{p}\Big{)}^{1/p}&\leq C\Big{(}\inf_{Q^{*}_{R,\delta}}(Mx_{n}-v)+R^{2}\sup|f|\Big{)}\\ &\leq C(R\inf_{F_{R,\delta}}\Big{(}M-\frac{v}{x_{n}})+R^{2}\sup|f|\Big{)}\\ &\leq C\Big{(}R\inf_{Q_{R/2,\delta}}(M-\frac{v}{x_{n}})+R^{2}\sup|f|\Big{)}\\ &\leq CR\Big{(}M-\sup_{Q_{R/2,\delta}}\frac{v}{x_{n}}+R\sup|f|\Big{)}.\end{split}$ (2.24)
@@
$\Big{(}R^{-n-2}\int_{Q^{\prime}_{R,\delta}}(v-mx_{n})^{p}\Big{)}^{1/p}\leq CR\Big{(}\inf_{Q_{R/2,\delta}}\frac{v}{x_{n}}-m+R\sup|f|\Big{)}.$ (2.25)
@@
$\operatorname{osc}\limits_{Q_{R,\delta}}\frac{v}{x_{n}}\leq C\Big{(}\frac{R}{R_{0}}\Big{)}^{\beta}\Big{(}\operatorname{osc}\limits_{Q_{R_{0},\delta}}\frac{v}{x_{n}}+R_{0}\sup|f|\Big{)}$ (2.26)
@@
where the constants $C$ and $\alpha$ depend only on $n,\epsilon,\Omega,c_{0},||\phi||_{\tilde{C}^{4}(\overline{Q}_{T})}$ and the uniform convexity of $\phi(\cdot,0)$.

### v2

@@ 2.1. Step 1: $C^{2}$ estimates on the boundary
$\mathcal{L}\mathcal{T}u=\mathcal{T}\psi$
for the linear operator $\mathcal{L}$ defined in (2.5). Thus, by (2.3), we have
(2.14) $\begin{split}|\mathcal{L}\mathcal{T}(u-\phi)|&\leq|\mathcal{T}\psi|+|\mathcal{L}\mathcal{T}\phi|\\ &\leq C\|D_{x}\psi\|_{L^{\infty}(Q_{T})}+|(u_{t}+\psi)u^{ij}\partial_{ij}\mathcal{T}\phi-\partial_{t}\mathcal{T}\phi|\\ &\leq\underline{c}\Big{(}1+\sum_{i=1}^{n}u^{ii}\Big{)}\end{split}$
for some positive constant $\underline{c}$ depending on $\Omega$, $\|D_{x}\psi\|_{L^{\infty}(Q_{T})}$, $||\phi||_{\tilde{C}^{3}(\overline{Q}_{T})}$ and the positive upper bound for $u_{t}+\psi$. Moreover, on $\Omega\times\{0\}$, $\mathcal{T}(u-\phi)=0$ and
(2.15) $|\mathcal{T}(u-\phi)|\leq C|x^{\prime}|^{2}|u_{n}-\phi_{n}|+\Big{|}\sum_{\beta<n}B_{\alpha\beta}x_{n}(u_{\beta}-\phi_{\beta})\Big{|}\leq\overline{c}x_{n},\quad\text{on }\partial\Omega\times[0,T],$
by the uniform convexity of $\Omega$ and the boundedness of $|Du|,|D\phi|$.
@@
Then $\mathcal{L}w=-2a(u_{t}+\psi)\sum_{i=1}^{n}u^{ii}$. Note that
$\sum_{i=1}^{n}u^{ii}\geq n\Big{(}\prod_{i=1}^{n}u^{ii}\Big{)}^{1/n}\geq n(\det D^{2}u)^{-1/n}\geq nC^{-1/n},$
where $C$ depends on the maximum of $u_{t}+\psi$. If we take $a$ large enough, then
@@
$u_{t}(0,t)+\psi(0,t)=\det D_{x^{\prime}x^{\prime}}u(0,t)u_{nn}(0,t)+P\big{(}D_{x^{\prime}x^{\prime}}u(0,t),D_{x^{\prime}x_{n}}u(0,t)\big{)},$
@@
(2.17) $|D^{2}u|\leq C\quad\text{ on }~{}~{}\partial_{p}Q_{T}.$
Since $u_{t}+\psi$ has a positive lower bound, then so does the smallest eigenvalue of $D^{2}u$ on the parabolic boundary according to (2.17). Thus estimate (2.9) is proved. ∎
@@ 2.2. Step 2: global $C^{2}$ estimates
(2.18) $-U_{t}-\frac{1}{\det D^{2}U}=-\psi(DU,t)\quad\text{ in }Q_{T}^{*}.$
Note that $Du(\Omega,t)$ is not necessarily convex. Here $Q_{T}^{*}$ is bounded and simply connected due to the smoothness of $u$ and formula (2.2). By estimates (2.3) and (2.9) above, we know that
(2.19) $C^{-1}\leq\det D^{2}U\leq C\quad\text{and}\quad||D^{2}U||_{L^{\infty}(\partial_{p}Q_{T}^{*})}\leq C.$
@@
Differentiating the equation (2.18), we have for $k,l=1,\cdots,n$,
$-U_{tk}+\frac{1}{\det D^{2}U}U^{ij}U_{ijk}=-\psi_{i}U_{ik}$
@@
$\begin{split}-U_{tkl}+\frac{1}{\det D^{2}U}U^{ij}U_{ijkl}-\frac{\big{(}U^{ij}U_{ijk}\big{)}\big{(}U^{ij}U_{ijl}\big{)}}{\det D^{2}U}-\frac{U_{ijk}U_{pql}U^{ip}U^{qj}}{\det D^{2}U}=-\psi_{ij}U_{ik}U_{jl}-\psi_{i}U_{ikl}.\end{split}$
If we define the linear operator $\tilde{L}=-\partial_{t}+\frac{1}{\det D^{2}U}U^{ij}\partial_{ij}+\psi_{i}\partial_{i}$, then
$\tilde{L}U_{kk}$ $=\frac{1}{\det D^{2}U}\Big{(}U^{ij}U_{ijk}\Big{)}^{2}+\frac{1}{\det D^{2}U}U_{ijk}U_{pqk}U^{ip}U^{qj}-\psi_{ij}U_{ik}U_{jk}.$
Since $U$ is convex and $\psi$ is concave, it holds that $U_{ijk}U_{pqk}U^{ip}U^{qj}\geq 0$ and $-\psi_{ij}U_{ik}U_{jk}\geq 0$. Therefore,
(2.20) $\tilde{L}U_{kk}\geq 0.$
By the comparison principle ([20, Theorem 2.4]) for linear parabolic equations, we obtain that
$U_{kk}\leq\sup_{\partial_{p}Q_{T}^{*}}U_{kk},$
@@
(2.21) $1/C\leq\lambda(D^{2}u)\leq C\quad\text{ on }~{}~{}\overline{Q}_{T},$
where $C$ is a positive constant depending on those universal quantities in Theorem 2.3.
@@ 2.3. Step 3: global $\tilde{C}^{2,\alpha}$ estimates
In this part, we first give the interior $\tilde{C}^{2,\alpha}$ regularity for $u$.
@@
Recall that $u\in\tilde{C}^{4}(\overline{Q}_{T})$ is a solution to equation (2.1) and is convex in $x$. Then, for any $Q\subset\subset Q_{T}$, there exists $\alpha\in(0,1)$ such that
$||u||_{\tilde{C}^{2,\alpha}(\overline{Q})}\leq C,$
where $\alpha,C$ are positive constants depending on those universal quantities in Theorem 2.3 and dist ($\text{dist\,}(Q,Q_{T})$.
Proof.
Recall that $U_{kk}$ satisfies (2.20) as $\psi(x,t)$ is concave with respect to $x$. By using the same method in [26, Theorem 4.1], it holds that $U\in\tilde{C}^{2,\alpha}(\Omega\times(0,T])$. Then by (2.21), we have the interior regularity for $u$ in $Q_{T}$. ∎
Recall that
$\mathcal{L}u_{t}=\psi_{t}\quad\text{ in }\ Q_{T},$
where the linear operator $\mathcal{L}=-\partial_{t}+\det D^{2}u\cdot u^{ij}\partial_{ij}$ is defined in (2.5). By estimate (2.21), the operator $\mathcal{L}$ is uniformly parabolic. Thanks to the Krylov-Safonov’s estimate for linear parabolic equations (see [8, Theorem 5.1] or [18]), we can obtain the global $\tilde{C}^{\alpha}$ estimate to $u_{t}$.
@@
for some positive constants $C$ and $\alpha$ depending on the uniformly parabolic coefficients of $\mathcal{L}$ and $\|\psi_{t}\|_{L^{n+1}(Q_{T})}$.
@@
The interior $\tilde{C}^{2,\alpha}$ regularity of $u$ can also be obtained by using the same method in [3, Proposition 5.3], once we get the global $C^{2}$ estimates and the $\tilde{C}^{\alpha}$ estimate for $u_{t}$.
We split the proof of the global $\tilde{C}^{2,\alpha}$ estimate into two parts. A short-time existence theorem gives the global $\tilde{C}^{\alpha}$ estimate for $D^{2}u$ on $\Omega\times(0,2\epsilon]$. As for the estimate on the rest part of $Q_{T}$, considering the linear parabolic equation satisfied by tangential derivatives of $u$, we use the approach of Theorem 9.31 in [10] to deduce the Hölder regularity of the gradient, which is independent of the moduli of continuity of the principal coefficients.
Let $Y_{0}=(y_{0},t_{0})$ be any point on $\partial\Omega\times[\epsilon,T]$ with some constant $\epsilon>0$. Without loss of generality, assume that $y_{0}=0$, and on a small neighborhood of 00, $\partial\Omega$ is the image of $y_{n}=\rho(y^{\prime})$ with $\rho(0)=0$ and $D\rho(0)=0$. Perform the transformation of coordinates $x=\Psi(y)$, and $\Psi:B_{\rho_{0}}(0)\cap\overline{\Omega}\to\overline{\mathbb{R}_{n}^{+}}$, with
@@
for a fixed $1\leq k<n$, where
$T_{i}:=\partial_{y_{i}}+\rho_{y_{i}}\partial_{y_{n}}\quad\text{ for }~{}i<n,\quad T_{n}=\partial_{y_{n}}.$
Take $G=B_{R_{0}}^{+}\times(t_{0}-R_{0}^{2},t_{0}]$, such that $B_{R_{0}}^{+}\subset\operatorname{Im}\Psi$ and $R_{0}^{2}<\epsilon$. Then by direct computations,
$\mathcal{L}(T_{k}u)$ $=-\partial_{t}(T_{k}u)+\det D^{2}u\cdot u^{ij}(T_{k}u)_{ij}$
$=T_{k}\psi+\det D^{2}u\cdot(u^{ij}\rho_{kij}u_{n}+u^{ij}\rho_{ki}u_{nj}+u^{ij}\rho_{kj}u_{ni})=:\overline{f}.$
Denote $P_{i}=\partial_{x_{i}}-\rho_{x_{i}}\partial_{x_{n}}$ for $1\leq i<n$ and $P_{n}=\partial_{x_{n}}$. Thus under $(x,t)$ coordinates,
$-\partial_{t}v+\det D^{2}u\cdot u^{ij}P_{i}P_{j}v=Lv-\det D^{2}u\cdot u^{ij}\rho_{ij}v_{n}.$
Here the linear operator $L=-\partial_{t}+a^{ij}\partial_{x_{i}x_{j}}$ with
$a^{ij}=\det D^{2}u\cdot(u^{ij}+u^{sl}\rho_{s}\rho_{l}\delta_{n}^{i}\delta_{n}^{j}-u^{sj}\rho_{s}\delta_{n}^{i}-u^{is}\rho_{s}\delta_{n}^{j}),$
where $\delta_{n}^{i}$ equals $1$ when $n=i$, otherwise 00, and the summation over $s,l$ is from $1$ to $n-1$. Note that $P_{j}v=\partial_{y_{j}}(T_{k}(u-\phi))$ and $P_{i}P_{j}v=\partial_{y_{i}y_{j}}(T_{k}(u-\phi))$ for $i,j=1,\cdots,n$.
Let
$f(x,t)=\Big{(}\det D^{2}u\cdot u^{ij}\rho_{ij}\partial_{y_{n}}(T_{k}(u-\phi))-\mathcal{L}(T_{k}\phi)+\overline{f}\Big{)}(\Psi^{-1}(x),t).$
@@
(2.22) {$-v_{t}+a^{ij}v_{ij}&=f\quad\text{ in }G,\\ v&=0\quad\text{ on }F,\end{aligned}\right.$
@@
(2.23) $\lambda I_{n}\leq(a^{ij})\leq\Lambda I_{n},$
if $\rho_{0}>0$ small, and
(2.24) $|f|\leq C_{0},\quad\quad|v|+|Dv|\leq C_{1}\quad\text{in}\quad\overline{G}$
for some positive constants $\lambda$, $\Lambda$, $C_{0}$ and $C_{1}$ depending on those universal quantities in Theorem 2.3.
@@
We will use Lemma 2.10 to prove an oscillation estimate for $\frac{v}{x_{n}}$, which gives the Hölder estimate for $v_{n}$ on $F$, the flat boundary. Combining this with the interior estimates, Lemma 2.9, and with the help of Lemma 2.11, we will derive a global $\tilde{C}^{1,\alpha}$ estimate for $v$ in $\frac{1}{2}G:=B_{R_{0}/2}^{+}\times(t_{0}-R_{0}^{2}/4,t_{0}]$. Recall the definition of $v$, then the global Hölder estimate of $D^{2}u$ on $\Omega\times(\epsilon,T]$ follows from equation (2.1).
@@
If we take $\sigma=r+\mu/\delta$, then $\sigma<2r$ and $X,X_{1}^{\prime}\in Q_{\sigma,\delta}\subset Q_{R_{0}/2,\delta}$. Consider the function $v(X)-\ell\cdot x$, then by Lemma 2.9 and Lemma 2.11,
(2.25) $\begin{split}|Dv(X_{1})-Dv(X_{1}^{\prime})|&\leq\sup_{Q(X_{1},\mu/2)}|Dv-\ell|\\ &\leq C\Big{(}\mu^{-1}\sup_{Q(X_{1},\mu)}|v(X)-\ell\cdot x|+\mu^{\alpha}\Big{)}.\end{split}$
@@
(2.26) $\operatorname{osc}\limits_{Q_{\sigma,\delta}}\frac{v(X)}{x_{n}}\leq C\sigma^{\alpha}.$
@@
$\leq Cr^{\alpha},$
where we use the fact that $\mu=\frac{1}{2}x_{1,n}<\delta r$ and $\sigma<2r$. Again by the claim,
$|Dv(X_{1}^{\prime})-Dv(0,t_{0})|\leq C\sigma^{\alpha}\leq Cr^{\alpha},$
then we have proved the theorem by the triangle inequality. For the rest part of this proof, we will show the claim (2.26).
@@
(2.27) $\inf_{F_{R,\delta}}\frac{v}{x_{n}}\leq 10\Big{(}\inf_{Q_{R/2,\delta}}\frac{v}{x_{n}}+R\sup|f|\Big{)}\quad\text{ for $R<1/4R_{0}$.}$
@@
Hence, $Lw\geq\sup|f|\geq Lv$ in $Q_{1,\delta}$, when $\delta=\delta(\lambda,\Lambda,c(n))>0$ is small enough. On $\{x_{n}=0\}\cap\partial_{p}Q_{1,\delta}$, $w=0=v$. On $\{x_{n}=\delta\}\cap\partial_{p}Q_{1,\delta}$, $w\leq\delta(1-|x^{\prime}|^{2})\leq\delta\leq v$ by (2.28). Additionally, on $\{|x^{\prime}|=1\}\cap\partial_{p}Q_{1,\delta}$ and $\{t=t_{0}-1\}\cap\partial_{p}Q_{1,\delta}$, $w<0\leq v$. Thus by maximal principle, we have $w\leq v$ in $Q_{1,\delta}$. Therefore, in $Q_{1/2,\delta}$,
@@
when $\delta<1/100$. Now the constant $\delta$ is determined, so does $r_{0}$.
Next, without the assumption (2.28), let
@@
(2.29) $\begin{split}\Big{(}R^{-n-2}\int_{Q^{\prime}_{R,\delta}}(Mx_{n}-v)^{p}\Big{)}^{1/p}&\leq C\Big{(}\inf_{Q^{*}_{R,\delta}}(Mx_{n}-v)+R^{2}\sup|f|\Big{)}\\ &\leq C(R\inf_{F_{R,\delta}}\Big{(}M-\frac{v}{x_{n}})+R^{2}\sup|f|\Big{)}\\ &\leq C\Big{(}R\inf_{Q_{R/2,\delta}}(M-\frac{v}{x_{n}})+R^{2}\sup|f|\Big{)}\\ &\leq CR\Big{(}M-\sup_{Q_{R/2,\delta}}\frac{v}{x_{n}}+R\sup|f|\Big{)}.\end{split}$
@@
(2.30) $\Big{(}R^{-n-2}\int_{Q^{\prime}_{R,\delta}}(v-mx_{n})^{p}\Big{)}^{1/p}\leq CR\Big{(}\inf_{Q_{R/2,\delta}}\frac{v}{x_{n}}-m+R\sup|f|\Big{)}.$
@@
(2.31) $\operatorname{osc}\limits_{Q_{R,\delta}}\frac{v}{x_{n}}\leq C\Big{(}\frac{R}{R_{0}}\Big{)}^{\beta}\Big{(}\operatorname{osc}\limits_{Q_{R_{0},\delta}}\frac{v}{x_{n}}+R_{0}\sup|f|\Big{)}$
@@
where the constants $C$ and $\alpha$ depend only on $\epsilon^{-1}$ and those universal quantities in Theorem 2.3.